OpenStudy (anonymous):

A car traveling 40 mi/h accelerates uniformly for 5.3 s, covering 551 ft in this time. What was its acceleration? Answer in units of ft/s 2 Ive gotten a=12.05 and 14.34 but neither of those are right.

OpenStudy (anonymous):

Did you convert your velocity from mi/h to ft/s correctly? There are 5280 ft/mi and 3600 s/h.

OpenStudy (anonymous):

40m/h = 58.7ft/s This is my way of doing this: v^2 = (v initial)^2 + 2as (58.7ft/s)^2 = 0 + 2a(551ft) 3442ft2/s2 = (1102ft)a => a = 3.12ft/s^2

OpenStudy (anonymous):

thats not the answer, i assumed we dont know final velocity, so i have initalv= 40mi/h, time=5.3seconds and d=551ft. I think i must use the equation d=vi*t+.5(a)(t)^2 but nothing i get is right.

OpenStudy (anonymous):

what is the right answer that's written in your textbook?

OpenStudy (anonymous):

its an online assignment so i dont know the right answer, it just lets me know if it is wrong. I know for sure thats the equation i use because the next question of the assignment asks what is the final velocity.

OpenStudy (anonymous):

can you try 17.1m/s2 to see if it's correct?

OpenStudy (anonymous):

i mean ft

OpenStudy (anonymous):

YES! thats right!, thank you!

OpenStudy (anonymous):

nice! Just thought you should know I use the formula d = di + vi*t + 1/2at^2 => 551 = 0 + 5.3*58.7 + 1/2a(5.3)^2 => 551 = 311 + 14a => a = 17.1 ft/s2

OpenStudy (anonymous):

to find the final velocity do i use the equation, a=vf-vi/t ?

OpenStudy (anonymous):

yes. Be careful. use 58.7ft/s instead of 40mi/h

OpenStudy (anonymous):

Thanks so much for all your help, have a good night! ... or day dependning on where you are!